# Explaining Text Classifiers with Counterfactual Representations

### v2

@@ Abstract
One well motivated explanation method for classifiers leverages counterfactuals which are hypothetical events identical to real observations in all aspects except for one categorical feature. Constructing such counterfactual poses specific challenges for texts, however, as some attribute values may not necessarily align with plausible real-world events. In this paper we propose a simple method for generating counterfactuals by intervening in the space of text representations which bypasses this limitation. We argue that our interventions are minimally disruptive and that they are theoretically sound as they align with counterfactuals as defined in Pearl’s causal inference framework. To validate our method, we conducted experiments first on a synthetic dataset and then on a realistic dataset of counterfactuals. This allows for a direct comparison between classifier predictions based on ground truth counterfactuals—obtained through explicit text interventions—and our counterfactuals, derived through interventions in the representation space. Eventually, we study a real world scenario where our counterfactuals can be leveraged both for explaining a classifier and for bias mitigation [^1] Code, data and supplementary material are available at: github.com/toinesayan/counterfactual-representations-for-explanation..
@@ 1 Introduction
Providing an explanation for the predictions made by a text classifier for a particular document is essential in situations where social bias could have detrimental consequences, for example when documents refer to individuals belonging to different social groups. One well motivated explanation method for classifiers leverages counterfactuals which are hypothetical individuals identical to real ones in all aspects except for one categorical feature that is being intervened on (Mothilal et al., 2020). Understanding how a classifier reacts to such fictitious individuals will indeed furnish an explanation for how it uses different pieces of information for its predictions. This approach to explanations has been well investigated in the context of social fairness (Kusner et al., 2017; Garg et al., 2019) but it obviously has a wider scope.
Creating counterfactual (CF) for text documents poses specific challenges when compared to producing CF for tabular data. One of these is related to the fact that it is often by no means obvious how to define a CF text for which the value of some text attribute is modified while everything else is kept unchanged. A number of recent works do, however, propose methods for constructing explicit counterfactuals at the text level in restricted contexts. Zeng et al. (2020) for instance propose to intervene on entities, separated from their context, to provide CF text samples that can be used for improving the generalization of a NER classifier under limited observational samples. Calderon et al. (2022) also intervenes on the text by replacing some domain-specific terms to create coherent counterfactuals also used for data augmentation purposes. Madaan et al. (2021) perform controlled text generation to enforce some user provided label.
@@
The aim of this work is not to achieve any kind of SOTA performance. Instead, we aim to demonstrate the usefulness of a simple and theoretically sound regression-based approach to generate counterfactual representations, which can serve as a strong baseline for a variety of tasks.
@@ 2 Related Work
Using Pearl’s causal inference framework for defining counterfactual fairness was pioneered in Kusner et al. (2017). This work motivates and formalizes the intuition that a classifier which is fair towards individuals belonging to different social groups should produce the same predictions for an actual individual and for a counterfactual individual belonging to a different group, other things being equal. It also stresses the importance of taking into account causal relationships between the variables that describe an individual when constructing fair classifiers. These relations are typically expressed with a DAG associated to the structural causal models (SCM) (Pearl, 2009; Peters et al., 2017) which describes the data generation process. One central observation is that fair predictors should only rely on variables that are non-descendant of protected variables in the causal graph. Finally, the authors describe an algorithm for training fair classifiers that uses a deconvolution approach. Our method for producing CFR for text representations could be used as a practical way of identifying which values of a sensitive text attribute imply a violation of the counterfactual fairness of a possibly biased classifier.
@@
Another line of work (Ravfogel et al., 2020; Xie et al., 2017; Elazar et al., 2021; Belrose et al., 2023; Feder et al., 2021), focuses on defining methods for concept erasure. The aim is still to build fair predictors that use data from which information on protected attributes has been “scrubbed”. However, achieving such erasure can be tricky due to the presence in the text of numerous factors correlated with the concept to be erased (De-Arteaga et al., 2019). To circumvent this difficulty, one possibility is to intervene on text representations rather than on texts themselves. Intervention methods on representations generally fall into two categories: adversarial methods and linear methods. The former rely on a gradient-reversal layer during training to produce representations that do not encode information about the protected attribute (Feder et al., 2021), but have been proven to fail at fully removing this information (Elazar and Goldberg, 2018). Focusing on linear methods, (Ravfogel et al., 2020; Elazar et al., 2021; Belrose et al., 2023; Shao et al., 2023) use projections that remove unwanted information from the representation space. In our work we also opt to intervene on representations using a closed form projector acting on the representation space as in Belrose et al. (2023). Two aspects of this approach are worth mentioning. First, there is no need to train a machine learning model, and therefore it requires minimal computational resources. Second, except for the erased information, this method preserves as much information as possible in a precise sense.
@@ 3.1 Background
Our aim is to define CFRs that can be used as reliable substitutes for genuine CF. Let us thus start by enumerating what we intuitively expect from a “good” CFR. For the sake of clarity, let’s assume that a sentence $s$ describes the emotional state of an individual and that it is represented by an embedding $X(s)\in\mathbb{R}^{d}$, obtained from a standard encoder like BERT. Suppose that $\widehat{Y}$ is a classifier for some discrete $Y$ like the emotional content conveyed by $s$, and that $Z(s)$ is a discrete protected attribute like the gender or the race of the individual referred to in $s$. We thus assume the causal graph is $Z\to X\to\widehat{Y}$  [^2] The causal relationships with $Y$ do not concern us because we focus on explaining predictions $\widehat{Y}$..
@@ 3.2 Making Minimal Interventions
To define a CFR as a minimally disruptive intervention we follow (Ravfogel et al., 2023; Belrose et al., 2023) which introduced the concept of linear guardedness that we now briefly review. It formalizes the intuition that only part of the information in $X$ is useful for predicting $Z$ with a linear predictor. Let thus $\eta(X)$ be a predictor for $Z$ and assume that the loss function $\ell(\eta,Z)$ is convex in its first argument, which is the case for the usual cross-entropy for instance. A representation $X^{\perp}$ is then said to linearly guard $Z$ (as a one-hot encoded variable for $k$ categories in $\{0,1\}^{k}$) if no linear predictor $\eta(X^{\perp})=\mathbf{W}X^{\perp}+\mathbf{b}$ is able to predict $Z$ better than a constant predictor $\eta(X^{\perp})\equiv\mathbf{b}$. More formally, $X^{\perp}$ as a function of the text $s$ should maximize the minimum expected loss over linear predictors:
@@
The linearity assumption is a strong one [^3] Non-linear concept erasure is still largely an open problem. Three lines of work have tackled it recently. Adversarial approaches (Feder et al., 2021), kernelized versions of linear erasure methods (Ravfogel et al., 2022) and more recently approaches leveraging rate distortion theory (Basu Roy Chowdhury et al., 2023). but this is the price to pay for having the useful equivalence (Belrose et al., 2023) that $X^{\perp}$ linearly guards $Z$ iff
@@
provided $\mathbf{P}$ satisfies $\mathrm{ker}(\mathbf{P})\supseteq\mathrm{im}(\boldsymbol{\Sigma}_{XZ})$. In words, the projector $\mathbf{P}$ should nullify the column space of the covariance matrix $\mathbf{\Sigma}_{XZ}:=\mathrm{Cov}[X,Z]$. In general $\mathbf{P}\neq\mathbf{P}^{\top}$ and thus the projection is oblique and is not unique [^4] In Belrose et al. (2023) this freedom is used to make $X^{\perp}=\mathbf{P}X$ as close as possible to the original $X$ in quadratic mean.. We simply use the orthogonal projector on $\mathrm{im}(\mathbf{\Sigma}_{XZ})^{\perp}$. Denoting $V^{\parallel}=\mathrm{im}(\boldsymbol{\Sigma}_{XZ})$, which has dimension $k-1$ (as $Z$ is one-hot), and $V^{\perp}$ its orthogonal complement in $\mathbb{R}^{d}$, we decompose the representation space $\mathbb{R}^{d}$ as $V^{\perp}\oplus V^{\parallel}$.
@@ 3.3 Relation with Pearl’s Framework
In our case $O=\{Z,X^{\perp},X^{\parallel},\widehat{Y}\}$. The DAG $G$ of the relevant SCM is shown in Figure 2. It expresses the fact that the prediction $\widehat{Y}$ depends on the protected variable $Z$ only through $X^{\parallel}$ and that $X^{\parallel}$ and $X^{\perp}$ are correlated as revealed by experiment and displayed in Figure 1.
@@
where $\boldsymbol{\Sigma}^{\parallel}(z),\mathbf{W}(z),\mathbf{b}(z)$ can be expressed as closed-form expressions involving the components of $\boldsymbol{\mu}(z)$ and $\boldsymbol{\Sigma}(z)$. An SCM which is compatible with the above can now be proposed by introducing appropriate noise variables $U_{Z},U^{\perp},U^{\parallel}$ and linear functions $f_{Z},f_{X^{\perp}},f_{X^{\parallel}}$ associated with the $Z,X^{\perp},X^{\parallel}$ nodes in $G$. The predictor $\widehat{Y}$ is identified with $f_{\widehat{Y}}$ and has no associated noise variable. As we shall argue below, the noise variable $U^{\parallel}:=(U^{\parallel}_{1},\dots,U^{\parallel}_{k})$ for $X^{\parallel}$ should have as many components as the number of values $Z$ can take which is $k$. We shall write $U^{\parallel}_{Z}$ to mean $U^{\parallel}_{z}$ when $Z=z$. Using the definitions for $\boldsymbol{\mu}^{\perp},\boldsymbol{\mu}^{\parallel}(x^{\perp},z)$ and $\boldsymbol{\Sigma}^{\perp\perp}$ introduced in (7) and (8) we then define the SCM which implements the distribution $P(O)$ and the causality relations defined by $G$ as
@@
Suppose now that we observe the evidence $e:=(Z(s),X^{\perp}(s),X^{\parallel}(s)):=(z,x^{\perp},x^{\parallel})$ for some text $s$. Abduction amounts to reading off the values $u^{\perp}$ and $(u^{\parallel}_{1},\dots,u^{\parallel}_{k})$ of the noise variables from (9). Once conditioned on the evidence, $U_{Z}$ and $U^{\perp}$ are obviously not random anymore, so neither are $Z$ and $X^{\perp}$. The $X^{\parallel}$ variable on the other hand remains stochastic because knowing that $Z=z$ only freezes $u^{\parallel}_{z}$ but not $u^{\parallel}_{z_{1}}$ for $z_{1}\neq z$. Next, the counterfactual distribution $P(X_{Z\leftarrow{z_{1}}}|E=e)$ is defined by acting on $Z$ to set its value to $z_{1}\neq z$ in the SCM (9) in which the only remaining noise is $U^{\parallel}_{z_{1}}$. The distribution $P(X^{\parallel}_{Z\leftarrow z_{1}}|E=e)$ is thus given by (8) by replacing $z$ by $z_{1}$. Its expectation $\boldsymbol{\mu}^{\parallel}(x^{\perp},z)$ is nothing but the $\parallel$ component of our CFR $x_{Z\leftarrow z}(x)$ as defined by a linear regression in (4). The $\perp$ component is deterministic $X^{\perp}_{Z\leftarrow z}=x^{\perp}$ in agreement with the second component in (4). In other words, our CFR is nothing but the expectation of a counterfactual as defined in Pearl’s causal inference framework for the SCM (9), thus justifying our claim.
@@ 4.1 Datasets
We introduce a new synthetic dataset, EEEC+, as an extension of the existing EEEC dataset (Feder et al., 2021). Both are well suited for evaluating the impact of protected attributes (the gender or perceived race of the individual referred to in a text) on downstream mood state classification. Compared with EEEC, besides increasing the diversity of templates, we also turned the binary race concept into a ternary one to extend the scope of evaluation of our CFR model. Information on the creation and structure of EEEC+ can be found in the Supplementary Material A.
@@
This realistic dataset is well suited for evaluating the causal effect of a concept on a sentiment classification task (Abraham et al., 2022). It includes both 2,299 original restaurant reviews from OpenTable and human-edited counterfactual reviews in which an aspect of the dining experience (food, service, ambiance or noise) was modified. The analysis of causal effects in CEBaB is facilitated through the creation of edit pairs. These are pairs of observations from the same edit set that differ in the value of one aspect. An edit set comprises an original observation and all observations edited from that original observation. Observations have been annotated with multiply-validated sentiment ratings at the aspect level (mostly Positive, Negative, or Unknown labels) and at the review level (1 star (very negative) to 5 stars (very positive)). In this article, we use CEBaB’s exclusive train set described in (Abraham et al., 2022) as training data. So we conducted the study on 5,117 observations distributed across three splits, with 1,755 training (34%), 1,673 validation (32%), and 1,689 test samples (33%).

### v3

@@ Abstract
One well motivated explanation method for classifiers leverages counterfactuals which are hypothetical events identical to real observations in all aspects except for one feature. Constructing such counterfactual poses specific challenges for texts, however, as some attribute values may not necessarily align with plausible real-world events. In this paper we propose a simple method for generating counterfactuals by intervening in the space of text representations which bypasses this limitation. We argue that our interventions are minimally disruptive and that they are theoretically sound as they align with counterfactuals as defined in Pearl’s causal inference framework. To validate our method, we conducted experiments first on a synthetic dataset and then on a realistic dataset of counterfactuals. This allows for a direct comparison between classifier predictions based on ground truth counterfactuals—obtained through explicit text interventions—and our counterfactuals, derived through interventions in the representation space. Eventually, we study a real world scenario where our counterfactuals can be leveraged both for explaining a classifier and for bias mitigation.
@@ 1 Introduction
Providing an explanation for the predictions made by a text classifier for a particular document is essential in situations where social bias could have detrimental consequences, for example when documents refer to individuals belonging to different social groups. One well motivated explanation method for classifiers leverages counterfactuals which are hypothetical individuals identical to real ones in all aspects except for one feature that is being intervened on (Mothilal et al., 2020). Understanding how a classifier reacts to such fictitious individuals will indeed furnish an explanation for how it uses different pieces of information for its predictions. This approach to explanations has been well investigated in the context of social fairness (Kusner et al., 2017; Garg et al., 2019) but it obviously has a wider scope.
Creating counterfactuals (CF) for text documents poses specific challenges when compared to producing CF for tabular data. One of these is related to the fact that it is often by no means obvious how to define a CF text for which the value of some text attribute is modified while everything else is kept unchanged. A number of recent works do, however, propose methods for constructing explicit counterfactuals at the text level in restricted contexts. Zeng et al. (2020) for instance propose to intervene on entities, separated from their context, to provide CF text samples that can be used for improving the generalization of a NER classifier under limited observational samples. Calderon et al. (2022) also intervene on the text by replacing some domain-specific terms to create coherent counterfactuals also used for data augmentation purposes. Madaan et al. (2021) perform controlled text generation to enforce some user provided label.
@@
The aim of this work is not to achieve any kind of SOTA performance. Instead, we aim to demonstrate the usefulness of a simple and theoretically sound regression-based approach to generate counterfactual representations, which can serve as a strong baseline for a variety of tasks. Code and data are available on github111github.com/ToineSayan/counterfactual-representations-for-explanation.
@@ 2 Related Work
Using Pearl’s causal inference framework for defining counterfactual fairness was pioneered in Kusner et al. (2017). This work motivates and formalizes the intuition that a classifier which is fair towards individuals belonging to different social groups should produce the same predictions for an actual individual and for a counterfactual individual belonging to a different group, other things being equal. It also stresses the importance of taking into account causal relationships between the variables that describe an individual when constructing fair classifiers. These relations are typically expressed with a DAG associated to the structural causal model (SCM) (Pearl, 2009; Peters et al., 2017) which describes the data generation process. One central observation is that fair predictors should only rely on variables that are non-descendant of protected variables in the causal graph. Finally, the authors describe an algorithm for training fair classifiers that uses a deconvolution approach. Our method for producing CFR for text representations could be used as a practical way of identifying which values of a sensitive text attribute imply a violation of the counterfactual fairness of a possibly biased classifier.
@@
Another line of work (Ravfogel et al., 2020; Xie et al., 2017; Elazar et al., 2021; Belrose et al., 2023; Feder et al., 2021), focuses on defining methods for concept erasure. The aim is still to build fair predictors that use data from which information on protected attributes has been “scrubbed”. However, achieving such erasure can be tricky due to the presence in the text of numerous factors correlated with the concept to be erased (De-Arteaga et al., 2019). To circumvent this difficulty, one possibility is to intervene on text representations rather than on texts themselves (barr2021counterfactual). Intervention methods on representations generally fall into two categories: adversarial methods and linear methods. The former rely on a gradient-reversal layer during training to produce representations that do not encode information about the protected attribute (Feder et al., 2021), but have been proven to fail at fully removing this information (Elazar and Goldberg, 2018). Focusing on linear methods, (Ravfogel et al., 2020; Elazar et al., 2021; Belrose et al., 2023; Shao et al., 2023) use projections that remove unwanted information from the representation space. In our work we also opt to intervene on representations using a closed form projector acting on the representation space as in Belrose et al. (2023). Two aspects of this approach are worth mentioning. First, there is no need to train a machine learning model, and therefore it requires minimal computational resources. Second, except for the erased information, this method preserves as much information as possible in a precise sense.
@@ 3.1 Background
Our aim is to define CFRs that can be used as reliable substitutes for genuine CF. Let us thus start by enumerating what we intuitively expect from a “good” CFR. For the sake of clarity, let’s assume that a sentence $s$ describes the emotional state of an individual and that it is represented by an embedding $X(s)\in\mathbb{R}^{d}$, obtained from a standard encoder like BERT. Suppose that $\widehat{Y}$ is a classifier for some discrete $Y$ like the emotional content conveyed by $s$, and that $Z(s)$ is a discrete protected attribute like the gender or the race of the individual referred to in $s$. We thus assume the causal graph is $Z\to X\to\widehat{Y}$. [^2] The causal relationships with $Y$ do not concern us because we focus on explaining predictions $\widehat{Y}$.
@@ 3.2 Making Minimal Interventions
To define a CFR as a minimally disruptive intervention we follow Ravfogel et al. (2023) which introduced the concept of linear guardedness that we now briefly review. It formalizes the intuition that only part of the information in $X$ is useful for predicting $Z$ with a linear predictor. Let thus $\eta(X)$ be a predictor for $Z$ and assume that the loss function $\ell(\eta,Z)$ is convex in its first argument, which is the case for the usual cross-entropy for instance. A representation $X^{\perp}$ is then said to linearly guard $Z$ (as a one-hot encoded variable for $k$ categories in $\{0,1\}^{k}$) if no linear predictor $\eta(X^{\perp})=\mathbf{W}X^{\perp}+\mathbf{b}$ is able to predict $Z$ better than a constant predictor $\eta(X^{\perp})\equiv\mathbf{b}$. More formally, $X^{\perp}$ as a function of the text $s$ should maximize the minimum expected loss over linear predictors:
@@
The linearity assumption is a strong one [^3] Non-linear concept erasure is still largely an open problem. Three lines of work have tackled it recently. Adversarial approaches (Feder et al., 2021), kernelized versions of linear erasure methods (Ravfogel et al., 2022) and more recently approaches leveraging rate distortion theory (Basu Roy Chowdhury et al., 2023). but this is the price to pay for having the useful equivalence in (Belrose et al., 2023) that $X^{\perp}$ linearly guards $Z$ iff
@@
provided $\mathbf{P}$ satisfies $\mathrm{ker}(\mathbf{P})\supseteq\mathrm{im}(\boldsymbol{\Sigma}_{XZ})$. In words, the projector $\mathbf{P}$ should nullify the column space of the covariance matrix $\mathbf{\Sigma}_{XZ}:=\mathrm{Cov}[X,Z]$. In general $\mathbf{P}\neq\mathbf{P}^{\top}$ and thus the projection is oblique and is not unique. [^4] In Belrose et al. (2023) this freedom is used to make $X^{\perp}=\mathbf{P}X$ as close as possible to the original $X$ in quadratic mean. We simply use the orthogonal projector on $\mathrm{im}(\mathbf{\Sigma}_{XZ})^{\perp}$. Denoting $V^{\parallel}=\mathrm{im}(\boldsymbol{\Sigma}_{XZ})$, which has dimension $k-1$ (as $Z$ is one-hot), and $V^{\perp}$ its orthogonal complement in $\mathbb{R}^{d}$, we decompose the representation space $\mathbb{R}^{d}$ as $V^{\perp}\oplus V^{\parallel}$.
@@ 3.3 Relation with Pearl’s Framework
In our case $O=\{Z,X^{\perp},X^{\parallel},\widehat{Y}\}$. The DAG $G$ of the relevant SCM is shown in Figure 2. It expresses the fact that the prediction $\widehat{Y}$ depends on the variable $Z$ only through $X^{\parallel}$ and that $X^{\parallel}$ and $X^{\perp}$ are correlated as revealed by experiment and displayed in Figure 1.
@@
where $\boldsymbol{\Sigma}^{\parallel}(z),\mathbf{W}(z),\mathbf{b}(z)$ can be expressed as closed-form expressions involving the components of $\boldsymbol{\mu}(z)$ and $\boldsymbol{\Sigma}(z)$. An SCM which is compatible with the above can now be proposed by introducing appropriate noise variables $U_{Z},U^{\perp},U^{\parallel}$ and linear functions $f_{Z},f_{X^{\perp}},f_{X^{\parallel}}$ associated with the $Z,X^{\perp},X^{\parallel}$ nodes in $G$. The predictor $\widehat{Y}$ is identified with $f_{\widehat{Y}}$ and has no associated noise variable. As we shall argue below, the noise variable $U^{\parallel}:=(U^{\parallel}_{1},\dots,U^{\parallel}_{k})$ for $X^{\parallel}$ should have as many components as the number of values $Z$ can take which is $k$. We shall write $U^{\parallel}(Z)$ to mean $U^{\parallel}_{z}$ when $Z=z$. Using the definitions for $\boldsymbol{\mu}^{\perp},\boldsymbol{\mu}^{\parallel}(x^{\perp},z)$ and $\boldsymbol{\Sigma}^{\perp\perp}$ introduced in (7) and (8) we then define the SCM which implements the distribution $P(O)$ and the causality relations defined by $G$ as
@@
$:=\boldsymbol{\mu}^{\parallel}(X^{\perp},Z)+\boldsymbol{\Sigma}^{\parallel}(Z)U^{\parallel}(Z)$
@@
Suppose now that we observe the evidence $e:=(Z(s),X^{\perp}(s),X^{\parallel}(s)):=(z,x^{\perp},x^{\parallel})$ for some text $s$. Abduction amounts to reading off the values $u^{\perp}$ and $(u^{\parallel}_{1},\dots,u^{\parallel}_{k})$ of the noise variables from (9). Once conditioned on the evidence, $U^{\parallel}_{z}$ and $U^{\perp}$ are obviously not random anymore, so neither are $Z$ and $X^{\perp}$. The $X^{\parallel}$ variable on the other hand remains stochastic because knowing that $Z=z$ only freezes $U^{\parallel}_{z}$ but not $U^{\parallel}_{z_{1}}$ for $z_{1}\neq z$. Next, the counterfactual distribution $P(X_{Z\leftarrow{z_{1}}}|E=e)$ is defined by acting on $Z$ to set its value to $z_{1}\neq z$ in the SCM (9) in which the only remaining noise is $U^{\parallel}_{z_{1}}$. The distribution $P(X^{\parallel}_{Z\leftarrow z_{1}}|E=e)$ is thus given by (8) by replacing $z$ by $z_{1}$. Its expectation $\boldsymbol{\mu}^{\parallel}(x^{\perp},z)$ is nothing but the $\parallel$ component of our CFR $x_{Z\leftarrow z}(x)$ as defined by a linear regression in (4). The $\perp$ component is deterministic $X^{\perp}_{Z\leftarrow z}=x^{\perp}$ in agreement with the second component in (4). In other words, our CFR is nothing but the expectation of a counterfactual as defined in Pearl’s causal inference framework for the SCM (9), thus justifying our claim.
@@ 4.1 Datasets
We introduce a new synthetic dataset, EEEC+, as an extension of the existing EEEC dataset (Feder et al., 2021). Both are well suited for evaluating the impact of protected attributes (the gender or perceived race of the individual referred to in a text) on downstream mood state classification. Compared with EEEC, besides increasing the diversity of templates, we also turned the binary race concept into a ternary one to extend the scope of evaluation of our CFR model. Information on the creation and structure of EEEC+ can be found in Supplementary Material A.
@@
This realistic dataset is well suited for evaluating the causal effect of a concept on a sentiment classification task (Abraham et al., 2022). It includes both 2,299 original restaurant reviews from OpenTable and human-edited counterfactual reviews in which an aspect of the dining experience (food, service, ambiance or noise) was modified. The analysis of causal effects in CEBaB is facilitated through the creation of edit pairs. These are pairs of observations from the same edit set that differ in the value of one aspect. An edit set comprises an original observation and all observations edited from that original observation. Observations have been annotated with multiply-validated sentiment ratings at the aspect level (mostly Positive, Negative, or Unknown labels) and at the review level (1 star (very negative) to 5 stars (very positive)). In this article, we use CEBaB’s exclusive train set described in Abraham et al. (2022) as training data. So we conducted the study on 5,117 observations distributed across three splits, with 1,755 training (34%), 1,673 validation (32%), and 1,689 test samples (33%).